Solve for 'r'in the following equation: 8!/(8-r)! = 336.
Answer by Alan3354(69443)

Solve for 'r'in the following equation: 8!/(8-r)! = 336.
----------
(8-r)! = 8!/336 = 120
120 = 5!
8-r = 5
r = 3